# rQUFguide.bib

### v1

@@ 5.1.1 Bid-Ask Spread and Price Moves:
Here $\mathbbm{1}_{(QD^{(\zeta)})}(t)$ is the indicator function for a QD event at side $\zeta$. As can be seen from the above, the spread process follows a jumpy point process methodology. Similarly the mid-price changes $dp(t)$ follow another jumpy point process. The only change is that while the spread is symmetrically changed by events in both the sides of the LOB, price changes happen anti-symmetrically. That is, an in-spread order at Ask side would decrease the mid-price but at the Bid side, it would increase the mid-price.
$dp(t)$ $=\sum_{\zeta}\frac{g(\zeta)}{2}\bigg{(}-\eta_{IS}dN_{LO_{\zeta_{IS}}}(t)+m^{(\zeta)}_{T}(t)\mathbbm{1}_{(QD^{(\zeta)})}(t)\big{(}dN_{CO_{\zeta_{T}}}(t)+dN_{MO_{\zeta}}(t)\big{)}\bigg{)}$ (27)
Here we define $g:\{ask,bid\}\rightarrow\{1,-1\}$ as a signum function over the two sides with $g(ask)=1$ and $g(bid)=-1$. It is clear that for large and medium tick stocks where price improvements of IS orders happen uniformly at 1 tick and the width of the top is 1 tick as well, we can see that the bid ask spread will change by 1 tick at an event. For small tick stocks we can expect multiple tick moves in the spread process governed both by the distribution of $\eta_{IS}$ as well as $m^{(\zeta)}_{T}(t)$.
@@ 5.1.2 Shape and Sparsity of the LOB:
$q^{(\zeta)}(x,t)$ $=Q^{(\zeta)}_{T}(t)\mathbbm{1}_{x=0.5s(t)}(t)+\frac{Q^{(\zeta)}_{D}(t)}{m^{(\zeta)}_{D}(t)}\mathbbm{1}_{x\geq 0.5s(t)+m^{(\zeta)}_{T}(t)}(t)$ (28)
@@
As can be seen from this, the shape of the book’s maxima will depend on the dynamics of $s(t)$ and $m^{(\zeta)}_{T}(t)$. Another noteworthy aspect from the above is that the instantaneous value of $q^{(\zeta)}(x,t)$ will be zero for $x\leq 0.5*s(t)$ which for small tick stocks means that the shape of the book will be increasing as we go deeper from the mid-price. In terms of sparsity, again the same fact that the book is empty from the mid-price to the top and again from the top of the book till the deeper level starts, sparsity should be replicated in the model as well.
5.2 Simulation Study:
@@
Using the model specified above, we simulate an order book with random Hawkes and LOB Parameters. In Figure 13(a), we show a sample of the simulated order book for a small tick configuration of the parameters. We select the Hawkes Parameters by taking a power law kernel of the form $\phi^{(.)}(t)=a(1+ct)^{-b}$ with values for parameters $a,b,c$ for the kernels and the exogenous intensities chosen from the calibrated results on AMZN.OQ from [jain2023hawkes]. The exact parameter specification is given in Appendix B. Note that this choice is purely driven by having a baseline LOB which makes logical sense (for eg, for a stable spread process, we require the rates of MOs and COs to approximately nullify the rates of LOs). We reflect the parameters of the Ask side to the Bid side to maintain model’s symmetry. For In-spread events, we choose $\alpha=95,\beta=0.6$ to be corresponding to a small-tick stock for this study. Following [jain2023hawkes], we choose geometric distributions for order sizes $\kappa_{(.)}$ with spikes at round numbers like 1, 10, and 100. Similarly for unseen queue sizes, we use the geometric distribution with spikes at 1, 10, 100, 500 and 1000. Next, for $\eta_{(.)}$, we choose the geometric distribution again (with no spikes) since empirical observations (Figure 7(a)) show that the distance between top and top + 1 levels seem to follow an exponential distribution.
We vary each of the parameters on a range of values and showcase the fact that this model is general enough to be able to simulate characteristics of all kinds of stocks regardless of their relative tick size.
5.2.1 Ergodicity:
We vary the initialization of the LOB and check for stationarity in the evolution of the LOB dynamics. In Figure 13(b), we vary the initial spread for a small-tick parameter configuration and observe that the spread quickly converges to the mean spread of the LOB process. In Figure 13(c), we vary the initial sparsity of the LOB model by varying the parameter $M0:=(M0_{T},M0_{D})$. We initialise the LOB’s $m_{T}(0),m_{D}(0)$ using the exponential distribution with parameters $M0_{T},M0_{D}$ respectively. We plot the the distribution of $m_{D}(t)$ and see that the distribution remains invariant of the choice of $M0$. These two figures reinforce our claim that the process formed from the dynamics illustrated above is ergodic.
(a) Example LOB Simulations
(b) Spread Timeseries: Varying $s(0)$ in the simulations
(c) Width of deep levels: Varying initial sparsity in the simulations
@@ 5.2.1 Ergodicity:
Figure 13: Simulation Study (Part 1): (a) showcases an example simulated coarse grained LOB - notice the spread is around 20-30 ticks and the top and top+1 levels have several empty levels as well. In (b) we do an ergodicity study for the spread time series by perturbing the initial spread of the simulation. Similarly in (c) we show the invariance of the distribution of $m_{D}(t)$ with various initial levels of sparsity.
5.2.2 Relative tick-size:
As can be seen in Equation 24, the in-spread events have two extra parameters $\alpha$ and $\beta$ which we respectively denote by AvgSpread and SpreadBeta. We postulate that controlling this parameter is vital for simulating LOBs of various relative tick sizes. In Figures 14(a) and 14(b), we can clearly see that varying this parameter gives us the characteristics of large, medium and small tick stocks respectively. However we note that absolute returns and sparsity is unchanged in these simulations (Figures 14(c) and 14(d)). To classify the simulations into small and large tick based on the spread distribution, we make use of the method of moments to calculate the shape parameter $k$ of the Weibull distribution. This is due to the fact that the Weibull distribution has a shape parameter $k$ which can control the distribution to look like an exponential distribution or like a log-normal. We plot the categories based on the classifier $k>1$ for small-tick assets and the rest for large-tick assets in Figure 15(a). We can clearly infer from the diagram that if we set $\beta$ to zero, i.e. make the in-spread orders’ intensity independent of the current spread, it is impossible to simulate a small tick order book. Finally we note that even though we see $k>1$ in all tests run with $\beta=0.9$, the mode of the spread observed in the simulations with $\alpha\leq 10$ ticks is still one-tick. This makes the simulated time series of spread very much alike to a large tick stock however the distribution of spread has a longer tail than the exponential distribution.
@@ 5.2.2 Relative tick-size:
Figure 14: Simulation Study (Part 2): In (a) we vary the IS Hawkes Parameters $\alpha,\beta$ (also known as AvgSpread, SpreadBeta resp.) and we can see the distribution of spread changes from one that looks similar to the exponential distribution to one that looks more like a log-normal distribution. The average shape of the LOB (in a coarse grained sense) is plotted in (b) for varying AvgSpread and SpreadBeta. We can clearly see the shift of the maximum liquidity from the top of the book (i.e. half tick away from mid price) to much deeper in the book as we decrease SpreadBeta and increase AvgSpread. In (c) and (d) we show that controlling these two parameters alone does not account for the returns distribution’s shape as well as the distribution of the state variable $m_{T}(t)$ to a very good extent.
@@
This leads us to the next crucial parameter for changing the relative tick-size of the simulated LOB - $\hat{\eta}_{(.)}$. This parameter controls the number of empty levels between two queues in the LOB and therefore is directly related to sparsity. We vary it from 0 to 1 where near one $\hat{\eta}$ would mean that, for example, the price improvements due to in-spread limit orders is always 1 tick and near zero $\hat{\eta}$ would correspond to price improvements happening randomly from 1 tick to spread-in-ticks. In Figures 16(a), 16(c), 16(d) and 16(b), we can see that varying this parameter leads to varying relative tick-size of the simulated LOBs even though the in-spread order intensity parameters ($\alpha,\beta$) are corresponding to that of a small-tick stock. Therefore we conclude that it is critical to calibrate both the $\alpha,\beta$ and $\hat{\eta}_{(.)}$ accurately. In terms of modelling the deeper part of the LOB, we show in Figures 16(e) and 16(f), that varying the sparsity levels can vary the variance of $m_{D}$ and the shape quite closely resembles that of the empirical data.
(a) Spread Distribution: Varying dynamic sparsity in the simulations
(b) Shape of the LOB: Varying dynamic sparsity in the simulations
(c) Absolute Returns: Varying dynamic sparsity in the simulations
(d) Distance between top and top + 1: Varying dynamic sparsity in the simulations
(e) $m_{D}$: Varying dynamic sparsity in the simulations
(f) $m_{D}$: AMZN Empirical
@@
Figure 16: Simulation Study (Part 3): In this set of figures we vary the dynamic sparsity parameters $\hat{\eta}_{(.)}$, while setting the $\alpha,\beta$ parameters to a small tick stock’s setting, and observe the simulated LOB dynamics. We can see that varying the dynamic sparsity can also enable us to vary the spread distribution’s mode, the shape of the book’s maxima, the returns distributions’ tail as well as the state variables’ ($m_{T}(t),m_{D}(t)$) distribution. For empirical shapes of $m_{T}(t)$, we refer to the figure 7(a) and we show an example small tick empirical distribution for $m_{D}(t)$ in (f).
5.3 Poisson Process:
@@
Keeping all else the same as in Section 5.1, we change the point process of the 12 events from a Hawkes Process to a Poisson Process with in-spread intensities again depend on the current spread as a power law (a similar comparison was performed in [jain2023hawkes]). We compare the stylized facts for a small tick stock parameter setting between the Hawkes model proposed above and the Poisson model.
@@
Figure 17: Hawkes Process vs Poisson Process: To compare with a simpler baseline, we make use of the Poisson Process (with similar handling of IS intensities as the Hawkes Process). While the spread distribution between the two look similar in shape, the returns distribution does not. Clearly the Hawkes Process simulations have a fatter tail than the Poisson Process. Secondly, we make use of an event ‘PriceChangeTime’ which is not explicitly modelled. Here we can see clearly that the Hawkes Process is able to create a bimodal distribution of inter-event times while the Poisson distribution remains multimodal. As noted in [jain2023hawkes], the former is more inline with empirical data as it captures the high frequency domain of the inter-event durations.
@@
As we can see in 17, with the methodology proposed above, one can replicate the spread distribution of a small tick stock even with a Poisson model of the events’ dynamics. However we note that the distribution shape and tail of returns in the Hawkes Process model is significantly different from that of the Poisson model. We also note that while the distribution of Price Change Time (defined as the time between too successive mid price changes) is bimodal in the Hawkes Process, it is unimodal in the Poisson Process. This quantity is of interest here since we do not explicitly model this event in the point process. As noted in [jain2023hawkes], the shape of returns and price change time distributions in the Hawkes Process method is closer to the empirical distribution shape than the Poisson Process method’s.
6 Assumptions and (Counter-) Evidence:
We make the following assumptions in this model and provide evidence for or against this assumption.

### v2

@@ 3.3.1 Bid-Ask Spread and Mid-Price Changes:
Here $\mathbbm{1}_{(QD^{(\zeta)})}(t)$ is the indicator function for a QD event at side $\zeta$. As can be seen from the above, the spread process follows a jumpy point process methodology. Similarly the mid-price changes $dp_{mid}(t)$ follow another jumpy point process. The only change is that while the spread is symmetrically changed by events in both the sides of the LOB, price changes happen anti-symmetrically. That is, an in-spread order at Ask side would decrease the mid-price but at the Bid side, it would increase the mid-price.
$dp_{mid}(t)$ $=\sum_{\zeta}\frac{z}{2}\bigg{(}-\eta_{IS}dN_{LO_{\zeta_{IS}}}(t)+m^{(\zeta)}_{T}(t)\mathbbm{1}_{(QD^{(\zeta)})}(t)\big{(}dN_{CO_{\zeta_{T}}}(t)+dN_{MO_{\zeta}}(t)\big{)}\bigg{)}$ (42)
It is clear that for large and medium tick stocks where price improvements of IS orders happen uniformly at one tick and the width of the top is one tick as well, we can see that the bid ask spread will change by one tick at an event. For small tick stocks we can expect multiple tick moves in the spread process governed both by the distribution of $\eta_{IS}$ as well as $m^{(\zeta)}_{T}(t)$.
@@ 3.3.2 Shape and Sparsity of the LOB:
$Q^{(\zeta)}(x,t)$ $=Q^{(\zeta)}_{T}(t)\mathbbm{1}_{x=\frac{1}{2}s(t)}(t)+\frac{Q^{(\zeta)}_{D}(t)}{m^{(\zeta)}_{D}(t)}\mathbbm{1}_{x\geq\frac{1}{2}s(t)+m^{(\zeta)}_{T}(t)}(t)$ (43)
@@
As can be seen from this, the shape of the book’s maxima will depend on the dynamics of $s(t)$ and $m^{(\zeta)}_{T}(t)$. Another noteworthy aspect from the above is that the instantaneous value of $Q^{(\zeta)}(x,t)$ will be zero for $x\leq\frac{s(t)}{2}$ which for small tick stocks means that the shape of the book will be increasing as we go deeper from the mid-price. In terms of sparsity, again the same fact that the book is empty from the mid-price to the top and again from the top of the book till the deeper level starts, sparsity should be replicated in the model as well.
4 Results from a Simulation Study
@@
Using the model specified above, we simulate an order book with random Hawkes and LOB Parameters. In Figure 12(a), we show a sample of the simulated order book for a small tick configuration of the parameters which is specified in the following. We select the Hawkes Parameters by taking a power law kernel of the form $\phi^{(.)}(t)=a(1+ct)^{-b}$ with values for parameters $a,b,c$ for the kernels and the exogenous intensities chosen from the calibrated results on AMZN.OQ from [jain2023hawkes]. The exact parameter specification is given in Appendix B. Note that this choice is purely driven by having a baseline LOB which makes logical sense (for eg, for a stable spread process, we require the rates of MOs and COs to approximately nullify the rates of LOs). We reflect the parameters of the Ask side to the Bid side to maintain model’s symmetry. For In-spread events, we choose $\alpha=0.95,\beta=0.6$ to be corresponding to a small-tick stock for this study. Following [jain2023hawkes], we choose geometric distributions for order sizes $\kappa_{(.)}$ with spikes at round numbers like 1, 10, and 100. Similarly for unseen queue sizes, we use the geometric distribution with spikes at 1, 10, 100, 500 and 1000. Next, for $\eta_{(.)}$, we choose the geometric distribution again (with no spikes) since empirical observations (Figure 7) show that the distance between top and top + 1 levels seem to follow a geometric distribution.
In the following sections we first perform an ergodicity study and then move on to identification of critical parameters in our model which decide whether the simulated LOB is a Large, Medium or Small tick LOB. Even though we start our analysis with a small tick configuration, we vary each of these parameters on a range of values to showcase the fact that this model is general enough to be able to simulate characteristics of all kinds of stocks regardless of their relative tick size. Note that this analysis is done by keeping all other parameters to be the same as presented in Appendix B and thus, does not represent a true calibration for Large tick assets. Even so, in this section we will show that controlling certain parameters can simulate a LOB very similar to a Large-tick LOB.
4.1 Ergodicity
(a) Example Simulated LOB Dynamics
(b) Example Empirical LOB Dynamics
(c) Spread Timeseries: Varying $s(0)$ in the simulations
(d) Width of deep levels: Varying initial sparsity in the simulations
@@ 4.1 Ergodicity
Figure 12: Simulation Study (Example LOB Dynamics and Ergodicity): (a) showcases an example simulated coarse grained LOB - notice the spread is around 20-30 ticks and the top and top+1 levels have several empty levels as well. (b) shows an empirical LOB (AMZN.OQ) with similar ‘fat-queues’. Notice the similarity in the two figures. In (c) we do an ergodicity study for the spread time series by perturbing the initial spread of the simulation. Similarly in (d) we show the invariance of the distribution of $m_{D}(t)$ with various initial levels of sparsity.
Since proving ergodicity analytically for Equations 41 and 42 under a Hawkes Counting Process for 12 types of events with various cross-interacting state variable dynamics is mathematically very difficult and beyond the scope of this paper, we resort to a simulation study to showcase the ergodic properties of the proposed model. In future work, we intend to perform an analysis for these equations in a simpler setting such as a Poisson Process with uniformly unit order sizes.
We vary the initialization of the LOB and check for stationarity in the evolution of the LOB dynamics. In Figure 12(c), we vary the initial spread for a small-tick parameter configuration and observe that the spread quickly converges to the mean spread of the LOB process. In Figure 12(d), we vary the initial sparsity of the LOB model by varying the parameter $M0:=(M0_{T},M0_{D})$. We initialise the LOB’s $m_{T}(0),m_{D}(0)$ using the exponential distribution with parameters $M0_{T},M0_{D}$ respectively. We plot the the distribution of $m_{D}(t)$ and see that the distribution remains invariant of the choice of $M0$. These two figures reinforce our hypothesis that the process formed from the dynamics illustrated above is ergodic.
@@
Figure 13: Simulation Study (IS Params): In (a) we vary the IS Hawkes Parameters $\alpha,\beta$ (also known as AvgSpread, SpreadBeta resp.) and we can see the distribution of spread changes from one that looks similar to the exponential distribution to one that looks more like a log-normal distribution. The average shape of the LOB (in a coarse grained sense) is plotted in (b) for varying AvgSpread and SpreadBeta. We can clearly see the shift of the maximum liquidity from the top of the book (i.e. half tick away from mid price) to much deeper in the book as we decrease SpreadBeta and increase AvgSpread. In (c) and (d) we show that controlling these two parameters alone does not account for the returns distribution’s shape as well as the distribution of the state variable $m_{T}(t)$ to a very good extent.
4.2 Controlling Relative tick-size : Controlling the Spread Distribution and Shape of the LOB
@@
As can be seen in Equation 39, the in-spread events have two extra parameters $\alpha$ and $\beta$. Since these two parameters control the events within the spread and therefore are a part of the process which control the width of the spread, we postulate that controlling this parameter is critical for simulating LOBs of various relative tick sizes. In Figures 13(a) and 13(b), we can clearly see that varying these parameters from $\alpha=0.01,\beta=0.95$ (i.e. a large-tick setting) to $\alpha=1.0,\beta=0.5$ (i.e. a small-tick setting) gives us the characteristics of large, medium and small tick stocks for the Bid-Ask Spread distribution and Shape of the LOB. However we note that absolute returns and sparsity is largely unchanged in these simulations (Figures 13(c) and 13(d)). This suggests that while $\alpha,\beta$ are critical parameters for spread distribution and LOB shape, they alone are insufficient to account for the stylized facts we are interested in. Regardless, to classify the simulations into small and large tick based on the spread distribution, we make use of the method of moments to calculate the Coefficient of Variation (i.e. $\sigma^{2}/\mu^{2}$ as described in Figure 2) of the Weibull distribution. This is due to the fact that the Weibull distribution has a shape parameter $k$ which can control the distribution to look like an exponential distribution or like a log-normal. This shape parameter depends solely on the CV of the data. We plot the categories based on the classifier $CV>1$ for small-tick assets and the rest for large-tick assets in Figure 14(a). Additionally, we report the maxima of the LOB shape (in ticks) in line with Section 2.3 for this grid over $\alpha,\beta$ in Figure 14(b). We can see the two phases of the simulated LOB in Figure 14(a): when $\alpha,\beta$ are high, the LOB simulated resembles a small tick LOB and when not, it is similar to a large-tick one. Additionally we can see the maxima of the LOB shape shift smoothly from two ticks to 20+ ticks in Figure 14(b) for the row with $\beta=0.7$, for example. Hence, using our model one can smoothly transition from a large tick LOB to a small tick LOB by varying the two IS parameters. Additionally, we can clearly infer from the diagram that if we set $\beta$ to zero, i.e. make the in-spread orders’ intensity independent of the current spread, it is impossible to simulate a small tick order book. This finding underscores the importance of a spread-dependent intensity for spread-changing events like the IS limit orders. Finally we note that even though we see $CV>1$ in all tests run with $\beta=0.9$, the mode of the spread observed in the simulations with $\alpha\leq 0.1\$$ is still one-tick. This makes the simulated time series of spread very much alike to a large tick stock however the distribution of spread has a longer tail than the exponential distribution. Therefore controlling $\beta$ alone might be insufficient and one needs to calibrate both the parameters in tandem.
4.3 Controlling Sparsity and Mid-Price Change Sizes
(a) Absolute Returns: Varying dynamic sparsity in the simulations
(b) Distance between top and top + 1: Varying dynamic sparsity in the simulations
(c) $m_{D}$: Varying dynamic sparsity in the simulations
(d) $m_{D}$: AMZN Empirical
@@ 4.3 Controlling Sparsity and Mid-Price Change Sizes
Figure 15: Simulation Study (Sparsity Params): In this set of figures we vary the dynamic sparsity parameters $\hat{\eta}_{(.)}$, while setting the $\alpha,\beta$ parameters to a small tick stock’s setting, and observe the simulated LOB dynamics. We can see that varying the dynamic sparsity can also enable us to vary the returns distributions’ tail as well as the state variables’ ($m_{T}(t),m_{D}(t)$) distribution. For empirical shapes of $m_{T}(t)$, we refer to the figure 7 and we show an example small tick empirical distribution for $m_{D}(t)$ in (f).
Continuing our discussion on how to create a more realistic LOB model for small tick stocks where the LOB is sparse and also exhibits multi-tick price move, we analyse the next crucial parameter for changing the relative tick-size of the simulated LOB - $\hat{\eta}_{(.)}$. This parameter controls the number of empty levels between two queues in the LOB and therefore is directly related to sparsity. We vary it from 0 to 1 where near one $\hat{\eta}$ would mean that, for example, the price improvements due to in-spread limit orders is always one tick and near zero $\hat{\eta}$ would correspond to price improvements happening randomly from one tick to spread-in-ticks. Unlike previous discussions, in 15(a) and 15(b), we show that this parameter can also affect the distribution of returns - note the higher variance with lower $\hat{\eta}_{(.)}$ akin to the small tick stocks shown in Figure 4 - as well as the sparsity of the LOB with the distribution of the top width variable $m_{T}(t)$ flattening as we decrease $\hat{\eta}_{(.)}$. In terms of modelling the deeper part of the LOB, we show in Figures 15(c) and 15(d), that varying the sparsity levels can vary the variance of $m_{D}$ and the shape quite closely resembles that of the empirical data. We can see that varying this parameter leads to varying relative tick-size of the simulated LOBs even though the in-spread order intensity parameters ($\alpha,\beta$) are corresponding to that of a small-tick stock i.e. $(0.95,0.6)$. We again stress that given the choice of $\beta=0$, it is impossible to simulate a small tick spread distribution even with high dynamic sparsity. Therefore we conclude that it is important to calibrate all of the critical parameters: $\alpha,\beta$ and $\hat{\eta}_{(.)}$ accurately.
4.4 Comparisons to a Poisson Process Model:
@@
In order to showcase the marginal gains in quality of fit we receive by using a Hawkes Process model for the events’ intensities rather than a simpler Poisson Process model, we create a baseline Poisson Process model with the same state variables and their dynamics. We study a few stylized facts and compare them to the Hawkes Model in the following. Keeping all else the same as in Section 3, we change the point process of the 12 events from a Hawkes Process to a Poisson Process with in-spread intensities again depend on the current spread as a power law (a similar comparison was performed in [jain2023hawkes]).
@@
Figure 16: Hawkes Process vs Poisson Process: To compare with a simpler baseline, we make use of the Poisson Process (with similar handling of IS intensities as the Hawkes Process). While the spread distribution between the two look similar in shape, the returns distribution does not. Clearly the Hawkes Process simulations have a fatter tail than the Poisson Process. Secondly, we make use of an event ‘PriceChangeTime’ which is not explicitly modelled. Here we can see clearly that the Hawkes Process is able to create a bimodal distribution of inter-event times while the Poisson distribution remains unimodal. As noted in [jain2023hawkes], the former is more inline with empirical data as it captures the high frequency domain of the inter-event durations.
@@
As we can see in Figure 16, with the methodology proposed above, one can replicate the spread distribution of a small tick stock even with a Poisson model of the events’ dynamics. However we note that the distribution shape and tail of returns in the Hawkes Process model is significantly different from that of the Poisson model. We also note that while the distribution of Price Change Time (defined as the time between too successive mid price changes) is bimodal in the Hawkes Process, it is unimodal in the Poisson Process. This quantity is of interest here since we do not explicitly model this event in the point process. As noted in [jain2023hawkes], the shape of returns and price change time distributions in the Hawkes Process method is closer to the empirical distribution shape than the Poisson Process method’s. The reason for the bimodal nature of the inter-event times distributions in the Hawkes model is due to the interaction terms since removing them gives us the low-frequency mode alone. This multimodality in the distributions of the Hawkes events’ times is an interesting observation. The exact reason for only one other mode in the distribution and the calculation of this mode’s value is reserved for future research.
5 Model Calibration and Assumption Checking:
5.1 Model Calibration:
Figure 17: $\hat{\eta}_{IS}$ Calibration: $\hat{\eta}_{IS}$ is the parameter which controls the sparsity of the LOB. Here we show three examples, one from each category of stocks, of empirical histograms of the sparsity levels. On a log-log scale it can be clearly seen that the distribution is similar to the geometric distribution. Inset we also show the calibrated parameters with this prior for all three stocks - we observe the calibrated parameters move from near one to near zero as we decrease the rel. tick size. This indicates an increase in sparsity with a decrease in rel. tick size.
Since the Hawkes Dynamics is assumed to be independent of the LOB state variables, and since the distributions of the order sizes, and order distances from top are unconditional on any state variable and stationary, we can decouple the calibration to a parametric estimation of the distributions and a non-parametric estimation of the Hawkes kernels. We suggest the methodology described in [jain2023hawkes], combining [kirchner2017estimation] and [bacry2016estimation], for the latter. The authors also provide visualization of the estimated kernel norms for INTC (Large Tick), AAPL (Medium Tick), TSLA and AMZN (Small Tick). For the former, we suggest making use of the Maximum Likelihood method with a prior that all these distributions are Geometric distributions. In Figure 17, we showcase the empirical $\eta_{IS}$ distributions (note the shape looks exactly like the Geometric distribution) and their calibrated parameter: $\hat{\eta}_{(.)}$ across three stocks: INTC (large tick), AAPL (medium tick) and GOOG (small tick). Optionally, following [jain2023hawkes, LuAbergel2018], one can add spikes in the probability distribution function to account for the traders’ preference of round numbers in the order sizes and queue sizes. For the partition functions, in the following section, we make use of a simple linear partition function - i.e. we assume that the shape of the book in the deeper LOB is flat. This is clearly a strong assumption (Section 2.3), and a better choice could be using the average shape of the book instead.
5.2 Assumptions and Their Empirical Evidence or Counter-Evidence
In the previous section we showcased the fact that the model we propose is general enough so that controlling three parameters $\alpha,\beta$ and $\hat{\eta}_{(.)}$ enables us to simulate an LOB with Large or Small tick characteristics. This underlines the universality of the model. However as with any model, there are a number of simplifying assumptions made in the model construction to foster a straightforward analysis, an easier calibration, and an understandable simulation study. In this section we state the assumptions made for the model and and provide evidence for or against these assumptions.